OpenStudy (anonymous):

factorise 6x^3 -15xy^2

OpenStudy (anonymous):

start with \[3x(2x^2-5y^2)\]

OpenStudy (hoblos):

3x(6x^2 - 5y^2) 3x( xsqrt2 - ysqrt5)( xsqrt2 + ysqrt5)

OpenStudy (anonymous):

then what hobolos said, if you are allowed square roots. get \[3x(\sqrt{2}x+\sqrt{5}y)(\sqrt{2}x-\sqrt{5}y)\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

good work